# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ A.1 Lipschitz Continuity Regarding Spatial Variable
Let $\mu(\mathrm{d}\mathbf{x})=\exp(-U(\mathbf{x}))\mathrm{d}\mathbf{x}$ be a probability measure on a compact set $\Omega\subseteq\mathbb{R}^{d}$ whose potential $U:\Omega\rightarrow\mathbb{R}$ is twice continuously differentiable. Then, for every $f\in C^{1}(\Omega)$,
$\mathrm{Var}_{\mu}(f)\geq\langle\mathbb{E}_{\mu}[\nabla_{\mathbf{x}}f],(\mathbb{E}_{\mu}[\nabla_{\mathbf{x}}^{2}U])^{-1}\mathbb{E}_{\mu}[\nabla_{\mathbf{x}}f]\rangle.$
Remark A.2.
[…]
$\mathrm{Cov}_{\mu}(\mathbf{x})\succcurlyeq\left(\mathbb{E}_{\mu}[\nabla_{\mathbf{x}}^{2}U(\mathbf{x})]\right)^{-1}$
with equality if $\mathbf{x}\sim\mathcal{N}(\mathbf{m},\mathbf{\Sigma})$ with $\mathbf{\Sigma}$ being positive definite.
Proof of Lemma 4.3.
Under Assumption 4.4, we have
$-\nabla^{2}\log p_{\mathbf{x},t,\sigma}(\mathbf{y})=\frac{1}{\sigma^{2}(1-t)}\mathbf{I}_{d^{*}}-\nabla^{2}\log{\widehat{p}}^{*}_{data}(\mathbf{y}).$
Therefore,
$\left(\frac{1}{\sigma^{2}(1-t)}-\alpha\right)\mathbf{I}_{d^{*}}\preccurlyeq-\nabla^{2}\log p_{\mathbf{x},t,\sigma}(\mathbf{y})\preccurlyeq\left(\frac{1}{\sigma^{2}(1-t)}+\alpha\right)\mathbf{I}_{d^{*}}.$
[…]
$\mathrm{Cov}_{p_{\mathbf{x},t,\sigma}}(\mathbf{y})\succcurlyeq\left(\frac{1}{\sigma^{2}(1-t)}+\alpha\right)^{-1}\mathbf{I}_{d^{*}}=\frac{\sigma^{2}(1-t)}{1+\alpha\sigma^{2}(1-t)}\mathbf{I}_{d^{*}}.$
Thus, we have
$\nabla^{2}\log q(\sqrt{1-t}\sigma,\mathbf{x})\succcurlyeq-\frac{\alpha}{1+\alpha\sigma^{2}(1-t)}\mathbf{I}_{d^{*}}\succcurlyeq-\alpha\mathbf{I}_{d^{*}}.$
By Brascamp-Lieb inequality, when $11-\frac{1}{\alpha\sigma^{2}}<t<$ 1,
$\mathrm{Cov}_{p_{\mathbf{x},t,\sigma}}(\mathbf{y})\preccurlyeq\left(\frac{1}{\sigma^{2}(1-t)}-\alpha\right)^{-1}\mathbf{I}_{d}=\frac{\sigma^{2}(1-t)}{1-\alpha\sigma^{2}(1-t)}\mathbf{I}_{d^{*}}.$
Then, we have
$\nabla^{2}\log q(\sqrt{1-t}\sigma,\mathbf{x})\preccurlyeq\frac{\alpha}{1-\alpha\sigma^{2}(1-t)}\mathbf{I}_{d^{*}}.$
Let $\frac{\alpha}{1-\alpha\sigma^{2}(1-t)}$ and $\frac{d^{*}}{\sigma^{4}(1-t)^{2}}-\frac{1}{\sigma^{2}(1-t)}$ be equal, we get
$d^{*}=\frac{\sigma^{2}(1-t)}{1-\alpha\sigma^{2}(1-t)}.$
Solving the above equation yields
$t=t^{*}:=1-\frac{d^{*}}{\sigma^{2}(1+\alpha d^{*})}>1-\frac{1}{\alpha\sigma^{2}}.$
Based on the discussion above, we obtain
$\nabla^{2}\log q(\sqrt{1-t}\sigma,\mathbf{x})\preccurlyeq g(t)\mathbf{I}_{d^{*}},$
where
$g(t):=\begin{cases}\frac{\alpha}{1-\alpha\sigma^{2}(1-t)},&t\in(0,t^{*}]\\ \frac{d^{*}}{\sigma^{4}(1-t)^{2}}-\frac{1}{\sigma^{2}(1-t)},&t\in(t^{*},1).\end{cases}$
By taking the derivative of $g(t)$, we can see that $g(t)$ is increasing on $(0,t^{*}]$ and decreasing on $(t^{*},1)$. Therefore, we have
$g(t)\leq g(t^{*})=\alpha(1+\alpha d^{*})>\alpha,$
which implies Lemma 4.3 holds.
A.2 Upper Bound for Derivatives with Respect to Time $t$
Lemma A.3.
The partial derivative of $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ with respect to time t satisfies:
$\partial_{t}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$
$=-\frac{1}{\sigma^{2}(1-t)^{2}}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{2\sigma^{4}(1-t)^{3}}\left(\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\left((\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}\right)-\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2}\right).$
Proof.
To ease notation, we define $\varphi_{t}(\mathbf{x})=\int_{\mathbf{y}}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|}{2\sigma^{2}(1-t)}}\widehat{p}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}$, which is the unnormalized version of $q(\sqrt{1-t}\sigma,\mathbf{x})$. Note that $\nabla\log\varphi_{t}(\mathbf{x})=\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$. By using the product rule of the derivatives, we have
$\partial_{t}\nabla\log\varphi_{t}(\mathbf{x})=\frac{\partial_{t}\nabla\varphi_{t}(\mathbf{x})}{\varphi_{t}(\mathbf{x})}-\frac{\partial_{t}\varphi_{t}(\mathbf{x})\nabla\varphi_{t}(\mathbf{x})}{(\varphi_{t}(\mathbf{x}))^{2}}.$
We first compute $\frac{\partial_{t}\nabla\varphi_{t}(\mathbf{x})}{\varphi_{t}(\mathbf{x})}$ as follows:
$\frac{\partial_{t}\nabla\varphi_{t}(\mathbf{x})}{\varphi_{t}(\mathbf{x})}$ $=\frac{\int_{\mathbf{y}}\left(-\frac{\mathbf{x}-\mathbf{y}}{\sigma^{2}(1-t)^{2}}+\frac{(\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{4}(1-t)^{3}}\right)e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}}{\varphi_{t}(\mathbf{x})}$
$=-\frac{1}{\sigma^{2}(1-t)^{2}}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})+\frac{1}{2\sigma^{4}(1-t)^{3}}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\left((\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}\right).$
By same calculus, we have
$\frac{\partial_{t}\varphi_{t}(\mathbf{x})}{\varphi_{t}(\mathbf{x})}=-\frac{1}{2\sigma^{2}(1-t)^{2}}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2},$
and
$\frac{\nabla\varphi_{t}(\mathbf{x})}{\varphi_{t}(\mathbf{x})}=-\frac{1}{\sigma^{2}(1-t)}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y}).$
Combining these three terms, we obtain
$\partial_{t}\nabla\log\varphi_{t}(\mathbf{x})$
$=-\frac{1}{\sigma^{2}(1-t)^{2}}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{2\sigma^{4}(1-t)^{3}}\left(\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\left((\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}\right)-\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2}\right).$
Lemma A.4.
${\rm{sup}}_{t\in[0,T]}{\rm{sup}}_{\mathbf{x}\in[-r,r]^{d}}\|{\partial_{t}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})}\|\leq\frac{\sqrt{d}(r+1)}{\sigma^{2}(1-T)^{2}}+\frac{d^{3/2}(r+1)^{3}}{\sigma^{4}(1-T)^{3}}.$
Proof.
From Lemma A.3, we have
$\|{\partial_{t}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})}\|$
$\leq\frac{1}{\sigma^{2}(1-t)^{2}}\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\|$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\frac{1}{2\sigma^{4}(1-t)^{3}}\left\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\left((\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}\right)-\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2}\right\|.$
By applying the Cauchy-Schwartz inequality, we obtain
$\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\|$ $\leq\left(\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2}\right)^{\frac{1}{2}}\leq\sqrt{d}(r+1),$
and
$\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\left((\mathbf{x}-\mathbf{y})\|{\mathbf{x}-\mathbf{y}}\|^{2}\right)\|\leq\left(\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{6}\right)^{\frac{1}{2}}\leq d^{3/2}(r+1)^{3},$
$\left\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\mathbb{E}_{p_{\mathbf{x},t,\sigma}}\|{\mathbf{x}-\mathbf{y}}\|^{2}\right\|\leq d(r+1)^{2}\|\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})\|\ \leq d^{3/2}(r+1)^{3}.$
The above inequalities imply
${\rm{sup}}_{t\in[0,T]}{\rm{sup}}_{\mathbf{x}\in[-r,r]^{d}}\|{\partial_{t}\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})}\|\leq\frac{\sqrt{d}(r+1)}{\sigma^{2}(1-T)^{2}}+\frac{d^{3/2}(r+1)^{3}}{\sigma^{4}(1-T)^{3}}.$
Appendix B Pre-training
In this section, we prove Theorem 4.1. We first decompose the excess risk error into the approximation error and the statistical error. Then, we bound these two terms separately. Finally, we balance these two terms to derive the error bound.
Proof of Theorem 4.1. We decompose the proof into the following four parts.
Error Decomposition. For each $\boldsymbol{E}_{\theta}\in\mathcal{E}$ and $\boldsymbol{D}_{\theta}\in\mathcal{D}$, we have
$\leq\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})+\sup_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})$
where the inequality is due to the fact that ^$\widehat{\mathcal{H}}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})\leq\widehat{\mathcal{H}}(\boldsymbol{E}_{\theta},\boldsymbol{D}_{\theta})$. Then taking infimum over $\boldsymbol{E}_{\theta}\in\mathcal{E}$ and $\boldsymbol{D}_{\theta}\in\mathcal{D}$ yields
[…]
The first and second terms in (10) are called the statistical error and approximation error respectively. Next, we bound these two terms separately.
Approximation Error. Since $\boldsymbol{E}^{*}$ and $\boldsymbol{D}^{*}$ are continuously differential on $[0,1]^{d}$, $[0,1]^{d^{*}}$ respectively, then there exist $B_{\boldsymbol{E}}$ and $B_{\boldsymbol{D}}$ such that
$\sup_{\mathbf{y}\in[0,1]^{d}}\|\boldsymbol{E}^{*}(\mathbf{y})\|\leq B_{\boldsymbol{E}},\leavevmode\nobreak\ \sup_{\mathbf{y}\in[0,1]^{d^{*}}}\|\boldsymbol{D}^{*}(\mathbf{y})\|\leq B_{\boldsymbol{D}}.$
For any given $10<\epsilon<$ 1, there exists a neural network $\boldsymbol{E}\in\mathcal{E}$ with configuration
@@ Appendix B Pre-training
$K_{\boldsymbol{E}}=\mathcal{O}(B_{\boldsymbol{E}}),\kappa_{\boldsymbol{E}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{E}}\}\right),$
such that
$\mathop{\sup}_{\mathbf{y}\in[0,1]^{d}}\|\boldsymbol{E}(\mathbf{y})-\boldsymbol{E}^{*}(\mathbf{y})\|_{\infty}\leq\epsilon.$
Moreover, the network $\boldsymbol{E}$ is Lipschitz continuous with Lipschitz constant $\gamma_{\boldsymbol{E}}=10d\xi_{\boldsymbol{E}}$, meaning that for any $\mathbf{y}_{1}$, $\mathbf{y}_{2}\in[0,1]^{d}$, it satisfies
@@
$\mathop{\sup}_{\mathbf{y}\in[0,1]^{d^{*}}}\|\boldsymbol{D}(\mathbf{y})-\boldsymbol{D}^{*}(\mathbf{y})\|_{\infty}\leq\epsilon,$
and the network $\boldsymbol{D}$ is Lipschitz continuous with Lipschitz constant $\gamma_{\boldsymbol{D}}=10d^{*}\xi_{\boldsymbol{D}}$, i.e., for any $\mathbf{y}_{1}$, $\mathbf{y}_{2}\in[0,1]^{d^{*}}$, it satisfies
@@
Therefore, for any $\mathbf{y}\in[0,1]^{d}$, we have
@@
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})\|$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E})(\mathbf{y})\|+\|(\boldsymbol{D}^{*}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})\|\right)$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})(\sqrt{d}\epsilon+\xi_{\boldsymbol{D}}\|\boldsymbol{E}(\mathbf{y})-\boldsymbol{E}^{*}(\mathbf{y})\|)$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})(\sqrt{d}\epsilon+\sqrt{d}\epsilon\xi_{\boldsymbol{D}})$
$=$ $\sqrt{d}\epsilon(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})(1+\xi_{\boldsymbol{D}}).$
@@
$\mathop{\inf}_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})$
$=$ $\mathop{\inf}_{\boldsymbol{E}\in\mathcal{E},\boldsymbol{D}\in\mathcal{D}}\int_{[0,1]^{d}}\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-\mathbf{y}\|^{2}-\|(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})-\mathbf{y}\|^{2}\right)\widetilde{p}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}$
$\leq$ $\sqrt{d}\epsilon(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+\sqrt{d})(1+\xi_{\boldsymbol{D}})$
$=$ $\mathcal{O}(\epsilon).$
Statistical Error. For convenience, we denote $\ell_{\boldsymbol{E},\boldsymbol{D}}(\mathbf{y}):=\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-\mathbf{y}\|^{2}$. We denote the $\delta$-covering of $\mathcal{D}$ with minimum cardinality $\mathcal{N}(\delta,\mathcal{D},\|\cdot\|_{L^{\infty}([0,1]^{d^{*}})})$ as $\mathcal{D}_{\delta}$ and the $\delta$-covering of $\mathcal{E}$ with minimum cardinality $\mathcal{N}(\delta,\mathcal{E},\|\cdot\|_{L^{\infty}[0,1]^{d}})$ as $\mathcal{E}_{\delta}$. $\mathcal{N}(\delta,\mathcal{D},\|\cdot\|_{L^{\infty}([0,1]^{d^{*}})})$ and $\mathcal{N}(\delta,\mathcal{E},\|\cdot\|_{L^{\infty}([0,1]^{d})})$ are bounded as follows:
$\log\mathcal{N}(\delta,\mathcal{D},\|\cdot\|_{L^{\infty}([0,1]^{d^{*}})})\lesssim J_{\boldsymbol{D}}L_{\boldsymbol{D}}\log\left(\frac{L_{\boldsymbol{D}}M_{\boldsymbol{D}}\kappa_{\boldsymbol{D}}}{\delta}\right),$
@@
$\log\mathcal{N}(\delta,\mathcal{E},\|\cdot\|_{L^{\infty}([0,1]^{d})})\lesssim J_{\boldsymbol{E}}L_{\boldsymbol{E}}\log\left(\frac{L_{\boldsymbol{E}}M_{\boldsymbol{E}}\kappa_{\boldsymbol{E}}}{\delta}\right).$
For any $\boldsymbol{E}\in\mathcal{E}$, $\boldsymbol{D}\in\mathcal{D}$, there exist $\boldsymbol{E}_{\delta}$, $\boldsymbol{D}_{\delta}$ such that
$\|\boldsymbol{E}(\mathbf{y})-\boldsymbol{E}_{\delta}(\mathbf{y})\|_{L^{\infty}([0,1]^{d})}\leq\delta,$
@@
$\|\boldsymbol{D}(\mathbf{y})-\boldsymbol{D}_{\delta}(\mathbf{y})\|_{L^{\infty}([0,1]^{d^{*}})}\leq\delta.$

### v2

@@ Appendix A Pre-training
where $\Psi_{\mathbf{v}}(\mathbf{y}^{\prime})$ is a partition of unity function which satisfies that
$\sum_{\mathbf{v}\in[N_{1}]^{d}}\Psi_{\mathbf{v}}(\mathbf{y}^{\prime})\equiv 1,\text{\leavevmode\nobreak\ on\leavevmode\nobreak\ }[0,1]^{d}.$
We choose $\Psi_{\mathbf{v}}$ as a product of coordinate-wise trapezoid functions, i.e.,
$\Psi_{\mathbf{v}}(\mathbf{y}^{\prime}):=\prod_{i=1}^{d}\psi\left(3N_{1}\left(y_{i}^{\prime}-\frac{v_{i}}{N_{1}}\right)\right),$
where $\psi$ is a trapezoid function, i.e.,
$\psi(x):=\left\{\begin{array}[]{ll}1,&|x|<1\\ 2-|x|,&|x|\in[1,2],\\ 0,&|x|>2.\end{array}\right.$
We claim that for $1\leq i\leq d^{*}$,
  • $[\overline{\boldsymbol{E}}(\mathbf{y}^{\prime})]_{i}$ is an approximation of $[\boldsymbol{E}^{*,\prime}(\mathbf{y}^{\prime})]_{i}$;
  • $[\overline{\boldsymbol{E}}(\mathbf{y}^{\prime})]_{i}$ can be implemented by a ReLU neural network $[\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime})]_{i}$ with small error.
Both claims are verified in [CLZZ20, Lemma 10], where we only need to substitute the Lipschitz constant $2\xi_{\boldsymbol{E}}$ into the error analysis. By concatenating $[\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime})]_{i}$, $1\leq i\leq d^{*}$ together, we construct $\boldsymbol{E}^{\prime}$. Given $\epsilon>0$, we can achieve
$\sup_{\mathbf{y}^{\prime}\in[0,1]^{d}}\|\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime})-\boldsymbol{E}^{*,\prime}(\mathbf{y}^{\prime})\|_{\infty}\leq\epsilon,$
where the neural network configuration satisfies
@@
$K_{\boldsymbol{E}}=\mathcal{O}(B_{\boldsymbol{E}}),\kappa_{\boldsymbol{E}}=\mathcal{O}\left(\max\{1,\xi_{\boldsymbol{E}}\}\right).$
Here, we set $e_{1}=\mathcal{O}\big{(}\frac{\epsilon}{\xi_{\boldsymbol{E}}}\big{)}$. Moreover, the neural network $\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime})$ is Lipschitz continuous. For any $\mathbf{y}^{\prime}_{1},\mathbf{y}^{\prime}_{2}\in[0,1]^{d}$, it holds that
$\|\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime}_{1})-\boldsymbol{E}^{\prime}(\mathbf{y}^{\prime}_{2})\|_{\infty}\leq 2\gamma_{\boldsymbol{E}}\|\mathbf{y}^{\prime}_{1}-\mathbf{y}^{\prime}_{2}\|,$
where $\gamma_{\boldsymbol{E}}:=10d\xi_{\boldsymbol{E}}$. Combining with the single input transformation layer, we denote
$\boldsymbol{E}(\mathbf{y}):=\boldsymbol{E}^{\prime}\left(\frac{\mathbf{y}+\mathbf{1}}{2}\right)\in\mathcal{E},$
then the configuration of $\boldsymbol{E}$ is the same as $\boldsymbol{E}^{\prime}$, and for any $\mathbf{y}_{1},\mathbf{y}_{2}\in[-1,1]^{d}$, it holds that
@@
$\mathop{\sup}_{\mathbf{y}\in[-1,1]^{d^{*}}}\|\boldsymbol{D}(\mathbf{y})-\boldsymbol{D}^{*}(\mathbf{y})\|_{\infty}\leq\epsilon,$
and the network $\boldsymbol{D}$ is Lipschitz continuous with Lipschitz constant $\gamma_{\boldsymbol{D}}:=10d^{*}\xi_{\boldsymbol{D}}$, i.e., for any $\mathbf{y}_{1}$, $\mathbf{y}_{2}\in[-1,1]^{d^{*}}$, it satisfies
@@
Now, we derive the upper bound of the approximation error. Since $\boldsymbol{E}^{*}$ and $\boldsymbol{D}^{*}$ are continuous differentiable on $[-1,1]^{d}$ and $[-1,1]^{d^{*}}$ respectively, then there exist two constants $B_{\boldsymbol{E}}$ and $B_{\boldsymbol{D}}$ such that
$\sup_{\mathbf{y}\in[-1,1]^{d}}\|\boldsymbol{E}(\mathbf{y})\|\leq B_{\boldsymbol{E}},\leavevmode\nobreak\ \sup_{\mathbf{y}\in[-1,1]^{d^{*}}}\|\boldsymbol{D}(\mathbf{y})\|\leq B_{\boldsymbol{D}}.$
Therefore, for any $\mathbf{y}\in[-1,1]^{d}$, we have
@@
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})\|$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E})(\mathbf{y})\|+\|(\boldsymbol{D}^{*}\circ\boldsymbol{E})(\mathbf{y})-(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})\|\right)$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})(\sqrt{d}\epsilon+\xi_{\boldsymbol{D}}\|\boldsymbol{E}(\mathbf{y})-\boldsymbol{E}^{*}(\mathbf{y})\|)$
$\leq$ $(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})(\sqrt{d}\epsilon+\sqrt{d}\epsilon\xi_{\boldsymbol{D}})$
$=$ $\sqrt{d}\epsilon(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})(1+\xi_{\boldsymbol{D}}).$
@@
$\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})=$ $\int_{[0,1]^{d}}\left(\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-\mathbf{y}\|^{2}-\|(\boldsymbol{D}^{*}\circ\boldsymbol{E}^{*})(\mathbf{y})-\mathbf{y}\|^{2}\right)\widetilde{p}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}$
$\leq$ $\sqrt{d}\epsilon(K_{\boldsymbol{D}}+B_{\boldsymbol{D}}+2\sqrt{d})(1+\xi_{\boldsymbol{D}})\lesssim\epsilon.$
The proof is complete. ∎
Proof of Theorem 4.2.
The proof can be divided into three parts. We first decompose the excess risk error into approximation error and statistical error. Then, we bound these two terms separately. Finally, we balance these two terms by choosing an appropriate $\epsilon$.
Error Decomposition. For each $\boldsymbol{E}_{\theta}\in\mathcal{E}_{0}$ and $\boldsymbol{D}_{\theta}\in\mathcal{D}_{0}$, we have
$\leq\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})+\sup_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}\widehat{\mathcal{H}}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E},\boldsymbol{D})$
where the inequality is due to the fact that ^$\widehat{\mathcal{H}}(\widehat{\boldsymbol{E}},\widehat{\boldsymbol{D}})\leq\widehat{\mathcal{H}}(\boldsymbol{E}_{\theta},\boldsymbol{D}_{\theta})$. Then taking infimum over $\boldsymbol{E}_{\theta}\in\mathcal{E}_{0}$ and $\boldsymbol{D}_{\theta}\in\mathcal{D}_{0}$ yields
[…]
The first and second terms in (11) are called the statistical error and approximation error respectively. By Lemma 4.1, the approximation error satisfies
$\inf_{\boldsymbol{E}\in\mathcal{E}_{0},\boldsymbol{D}\in\mathcal{D}_{0}}\mathcal{H}(\boldsymbol{E},\boldsymbol{D})-\mathcal{H}(\boldsymbol{E}^{*},\boldsymbol{D}^{*})\lesssim\epsilon.$
Statistical Error. For convenience, we denote $\ell_{\boldsymbol{E},\boldsymbol{D}}(\mathbf{y}):=\|(\boldsymbol{D}\circ\boldsymbol{E})(\mathbf{y})-\mathbf{y}\|^{2}$. We denote the $\delta$-covering of $\mathcal{D}_{0}$ with minimum cardinality $\mathcal{N}(\delta,\mathcal{D}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d^{*}})})$ as $\mathcal{D}_{\delta}$ and the $\delta$-covering of $\mathcal{E}_{0}$ with minimum cardinality $\mathcal{N}(\delta,\mathcal{E}_{0},\|\cdot\|_{L^{\infty}[-1,1]^{d}})$ as $\mathcal{E}_{\delta}$. $\mathcal{N}(\delta,\mathcal{D}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d^{*}})})$ and $\mathcal{N}(\delta,\mathcal{E}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d})})$ are bounded as follows:
$\log\mathcal{N}(\delta,\mathcal{D}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d^{*}})})\lesssim J_{\boldsymbol{D}}L_{\boldsymbol{D}}\log\left(\frac{L_{\boldsymbol{D}}M_{\boldsymbol{D}}\kappa_{\boldsymbol{D}}}{\delta}\right),$
@@
$\log\mathcal{N}(\delta,\mathcal{E}_{0},\|\cdot\|_{L^{\infty}([-1,1]^{d})})\lesssim J_{\boldsymbol{E}}L_{\boldsymbol{E}}\log\left(\frac{L_{\boldsymbol{E}}M_{\boldsymbol{E}}\kappa_{\boldsymbol{E}}}{\delta}\right).$
For any $\boldsymbol{E}\in\mathcal{E}_{0}$, $\boldsymbol{D}\in\mathcal{D}_{0}$, there exist $\boldsymbol{E}_{\delta}\in\mathcal{E}_{\delta}$, $\boldsymbol{D}_{\delta}\in\mathcal{D}_{\delta}$ such that
$\|\boldsymbol{E}(\mathbf{y})-\boldsymbol{E}_{\delta}(\mathbf{y})\|_{L^{\infty}([-1,1]^{d})}\leq\delta,$
@@
$\|\boldsymbol{D}(\mathbf{y})-\boldsymbol{D}_{\delta}(\mathbf{y})\|_{L^{\infty}([-1,1]^{d^{*}})}\leq\delta.$